the sum of 2 integers is 36. the difference is 4. find the integers

Answer by swincher4391(1107)

x+y = 36 x-y = 4
Use elimination method:

x+y = 36
x-y + 4
2x = 40
x = 20
20+y = 36
y = 16
20,16